A Leisure Time Cruises riverboat travels 46 km downstream in 2 hours. It travels 51 km upstream in 3 hours. Find the speed of the boat and the speed of the water.
Answer by mananth(16949)

A Leisure Time Cruises riverboat travels 46 km downstream in 2 hours. It travels 51 km upstream in 3 hours. Find the speed of the boat and the speed of the water.
rate of boat = b
rate of current = c
...
upstream speed = b-c mph
downstream speed = b+c mph
..
t= d/r
Upstream
51/(b-c)=3
multiply by (b-c)
51=3(b-c)
/3
17 = b-c
b-c = 17............................1
...
downstream
46/(b+c)=2
multiply by (b+c)
46= 2(b+c)
/2
b+c=23..............................2
...
add 1 & 2
2b=17+23
2b=40
/2
b=40/2
b=20 mph
...
b-c=17
20-c=17
c=3 mph
...
m.ananth@hotmail.ca